6. An acid-base reaction following the subsequent reaction occurs H2A + 2B- ⇌ 2HB + A2- a) What are the acid, base, conjugate acid, and conjugate base in this reaction? b) If H2A is a strong acid that fully deprotonates its hydrogens upon entering water, and B- is a weak base with a pKb value of 3.57 at room temp, what is the pH of a solution where we add 3.00 mL of 1.00M H2A to a 15.00mL solution of 0.500 M B- if the reaction is conducted at room temperature? c) Is the solution created in part b considered a buffer? How do we know? d) How much H2A would we need to add to reach the equivalence point?
Answer #1

a)    acid = H2A

        base = B-

       conjugate acid = HB

      conjugate base = A2-

b)

millimoles of H+ = 2 x 3 x 1.0 = 6

millimoles of B = 15 x 0.5 = 7.5

B    +    H+ -----------------> BH

7.5        6                             0

-6     -6                               +6

1.5     0                               6.

pOH = pKb + log (BH/B)

pOH = 3.57 + log (6/1.5)

pOH = 4.17

pH + pOH = 14

pH = 9.83

c)

yes it is buffer . because weak base B and its salt BH remains in the solution so it should be buffer.

d)

M1 V1 / n1 = M2 V2 / n2

1.0 x V1 / 1 = 0.5 x 15 / 2

V1 = 3.75 mL

volume of H2A needed for equilvalence point = 3.75 mL